### v1

@@ 5.5 Abraham-Lorentz Equations
$\frac{2}{3}\frac{e^{2}}{c^{3}}\frac{d\mathbf{a}}{d\tau}$ $=m\mathbf{a}-\frac{e}{c}\mathbf{v}\times\mathbf{B}(\mathbf{x}),$ (97a)
$\frac{d\mathbf{v}}{d\tau}$ $=\mathbf{a},$ (97b)
$\frac{d\mathbf{x}}{d\tau}$ $=\mathbf{v},$ (97c)
@@
$\left(\frac{r_{0}}{cT}\right)\frac{2}{3}\frac{d\mathbf{a}}{d\tau}$ $=\mathbf{a}-\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x}),$ (98a)
$\frac{d\mathbf{v}}{d\tau}$ $=(\omega_{c}T)\mathbf{a},$ (98b)
$\frac{d\mathbf{x}}{d\tau}$ $=\mathbf{v},$ (98c)
where we have introduced the cyclotron frequency $\omega_{c}=|e|B_{0}/(mc)$, the classical electron radius $r_{0}=e^{2}/(mc)^{2}$, and the sign of the charge $\zeta\in\{-1,+1\}$, where $\zeta=-1$ corresponds to electrons and $\zeta=+1$ to positrons. The dimensionless parameters
$\gamma_{R}=\frac{r_{0}}{cT},\qquad\gamma_{B}=\frac{1}{\omega_{c}T},$ (99)
@@
$\frac{d\mathbf{a}}{dt}$ $=\mathbf{a}-\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x}),$ (100a)
$\frac{d\mathbf{v}}{dt}$ $=\epsilon\frac{2}{3}\mathbf{a},$ (100b)
$\frac{d\mathbf{x}}{dt}$ $=\epsilon\frac{2}{3}\mathbf{v},$ (100c)
We will refer to (5.5) as the forward-time Abraham-Lorentz system. A normally-stable fast-slow system can be obtained by reversing time using $t_{-}=-t$,
$\frac{d\mathbf{a}}{dt_{-}}$ $=-\mathbf{a}+\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x}),$ (101a)
$\frac{d\mathbf{v}}{dt_{-}}$ $=-\epsilon\frac{2}{3}\mathbf{a},$ (101b)
$\frac{d\mathbf{x}}{dt_{-}}$ $=-\epsilon\frac{2}{3}\mathbf{v}.$ (101c)
We refer to (5.5) as the reverse-time Abraham-Lorentz system. In both systems, $\mathbf{a}$ is the fast variable and $\mathbf{v},\mathbf{x}$ are the slow variables. In reverse time, the acceleration, $\mathbf{a}$ will be attracted to the slow manifold, given by
$\mathbf{a}^{*}(\mathbf{x},\mathbf{v})=\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x})+\epsilon\frac{3}{2}\mathbf{v}\times\mathbf{B}(\mathbf{x})\times\mathbf{B}(\mathbf{x})+\mathcal{O}(\epsilon^{2}).$ (102)
@@
It is of interest to obtain forward-time trajectories of electrons along the slow manifold. A standard numerical integration of (5.5) will result in unstable trajectories on the fast time scale since any numerical error will push trajectories off of the slow-manifold, resulting in exponential growth of the acceleration. We instead apply FSNNs to learn the Fenichel normal form of the system in reverse time. We then obtain forward-time trajectories along the learned slow manifold by performing an integration of the slow variables along $y=0$ using (3.4). This test is therefore a more practical application of a data-driven closure.
We train a FSNN to learn the Abraham-Lorentz equations from data. The trajectory dataset, $\mathcal{D}$, consists of $10^{4}$ trajectories, corresponding to initial conditions randomly generated in the region, $\mathcal{R}$, specified by
@@
where $r=\sqrt{x_{1}^{2}+x_{2}^{2}}$ and $z=x_{3}$. $\epsilon$ is randomly generated log-uniformly in the range $[10^{-5},10^{-2}]$. The trajectories are integrated in reverse time using equations (5.5) using an adaptive RK45 solver set to an error tolerance of $10^{-8}$. The first 40 time steps batched in sequences of length 10 are used as fast-scale training data. These trajectories are also interpolated over a slower time scale with $\Delta t=0.025/\epsilon$ and $10^{4}$ sequences of length 10 are chosen at random for slow-scale training data. The manifold dataset, $\mathcal{M}$ consists of $10^{4}$ examples of data pairs $((\mathbf{x},\mathbf{v},\mathbf{a}),\epsilon=0)$ on the graph $\mathbf{a}=\mathbf{v}\times\mathbf{B}(\mathbf{x})$, where $\mathbf{x},\mathbf{v}$ are randomly generated in the region $\mathcal{R}$.
@@
$B_{r}$ $=-\frac{z}{q(r,z)r}B_{0},\qquad B_{\phi}=\frac{R_{0}}{r}B_{0},\qquad B_{z}=\frac{(r-R_{0})}{q(r,z)r}B_{0},$ (103)
@@
$\mathbf{B}=\frac{1}{R}\nabla\psi\times\mathbf{e}_{\phi}+\frac{F_{0}}{R}\mathbf{e}_{\phi}.$ (104)
@@
A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 50. The function $g$ on the right-hand-side of the neural ODE uses a single-layer feedforward neural network with hidden dimension 100 and is added to a bilinear form network with rank 1. The remaining feedforward neural networks on the right-hand-side of the neural ODE system each consist of 50 layers and the bilinear form has a rank of 4.
Figure 9 shows comparisons of the solutions in $r{\rm-}z$ space between the trained network and trajectories computed using RK45 on the reverse-time system. Using the reverse-time trajectories, endpoints were supplied as initial conditions to the trained network to predict forward in time. For both the reverse-time predictions (red arrows) and forward-time predictions (green arrows), the the trained model agrees well with the solution computed numerically. The corresponding $L_{2}$ errors are shown for position, velocity, and acceleration in Figure 10, demonstrating that errors are low when integrating on the slow time scale.
Table 2: Comparison of true eigenvalues to the eigenvalues of the trained model for the Abraham-Lorentz problem.
@@
Figure 11 shows comparisons of orbits between the trained model and orbits obtained when integrating the system by substituting the first two terms of the asymptotic expansion in (102). Note again that when integrating along the slow manifold, only the slow variable is evolved using the transformation $h$ and the closure $y=0$. The trained model is able to qualitatively capture the gyro-motion behavior while also accurately capturing the guiding center dynamics over the slow time scale. Despite training using a limited number of trajectories over a relatively short time frame (2 or less gyration periods) due to the slow time scale instability of the reverse-time system, the trained network is highly capable of predicting long-time dynamics in forward-time for many more gyration cycles. We remind the reader that standard numerical methods cannot integrate the forward-time Abraham-Lorentz system along the slow manifold without experiencing an exponential growth instability. Direct integration of the asymptotic expansion removes the instability, but introduces an error dependent on $\epsilon$ due to the asymptotic approximation. This result highlights the unique capability of the proposed network.
@@
Figure 11: Comparison of orbits of the forward-time Abraham-Lorentz system (5.5) along the slow manifold for the first two terms of the asymptotic expansion in (102) (left) and the trained model (right) for two different initial conditions and choices of $\epsilon$.
@@ 6 Conclusions
In this paper, we introduced the fast-slow neural network (FSNN), which is a data-driven approach suitable for learning singularly perturbed dynamical systems where the fast-scale dynamics are dissipative. Our method enforces the existence of a trainable, attracting invariant slow manifold as a hard constraint. Invertible neural networks and neural ODEs are key components of the FSNN. We introduce bi-Lipschitz affine transformation (bLAT) layers for enforcing regularity of the invertible neural network. We also develop a negative Schur form network which parameterizes matrices with eigenvalues to the left of the imaginary axis. This in addition to the use of an additive L-stable diagonally-implicit Runge Kutta scheme for integration are key pieces in ensuring that the fast-scale dynamics are stable. The Schur form network also enables the use of back-substitution to efficiently handle the implicit solve. Analytical representation of the slow manifold enables efficient integration on the slow time scale. We demonstrate the FSNN on many examples that exhibit two timescales, including the Grad moment system from hydrodynamics, two-scale Lorentz96 equations for modeling atmospheric dynamics, and Abraham-Lorentz dynamics modeling radiation reaction of electrons in a magnetic field.
The proposed network can be used to discover a full-order model with long-term accuracy near a slow manifold when only short duration trajectories are available. The proposed network is also applicable to discover a reduced-order model as a unique closure discovery procedure.

### v2

@@ 5.5 Abraham-Lorentz Equations
(96a) $\frac{2}{3}\frac{e^{2}}{c^{3}}\frac{d\mathbf{a}}{d\tau}$ $=m\mathbf{a}-\frac{e}{c}\mathbf{v}\times\mathbf{B}(\mathbf{x}),$
(96b) $\frac{d\mathbf{v}}{d\tau}$ $=\mathbf{a},$
(96c) $\frac{d\mathbf{x}}{d\tau}$ $=\mathbf{v},$
@@
(97a) $\left(\frac{r_{0}}{cT}\right)\frac{2}{3}\frac{d\mathbf{a}}{d\tau}$ $=\mathbf{a}-\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x}),$
(97b) $\frac{d\mathbf{v}}{d\tau}$ $=(\omega_{c}T)\mathbf{a},$
(97c) $\frac{d\mathbf{x}}{d\tau}$ $=\mathbf{v},$
where we have introduced the cyclotron frequency $\omega_{c}=|e|B_{0}/(mc)$, the classical electron radius $r_{0}=e^{2}/(mc^{2})$, and the sign of the charge $\zeta\in\{-1,+1\}$, where $\zeta=-1$ corresponds to electrons and $\zeta=+1$ to positrons. The dimensionless parameters
(98) $\gamma_{R}=\frac{r_{0}}{cT},\qquad\gamma_{B}=\frac{1}{\omega_{c}T},$
@@
(99a) $\frac{d\mathbf{a}}{dt}$ $=\mathbf{a}-\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x}),$
(99b) $\frac{d\mathbf{v}}{dt}$ $=\epsilon\frac{2}{3}\mathbf{a},$
(99c) $\frac{d\mathbf{x}}{dt}$ $=\epsilon\frac{2}{3}\mathbf{v},$
We will refer to (99) as the forward-time Abraham-Lorentz system. A normally-stable fast-slow system can be obtained by reversing time using $t_{-}=-t$,
(100a) $\frac{d\mathbf{a}}{dt_{-}}$ $=-\mathbf{a}+\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x}),$
(100b) $\frac{d\mathbf{v}}{dt_{-}}$ $=-\epsilon\frac{2}{3}\mathbf{a},$
(100c) $\frac{d\mathbf{x}}{dt_{-}}$ $=-\epsilon\frac{2}{3}\mathbf{v}.$
We refer to (100) as the reverse-time Abraham-Lorentz system. In both systems, $\mathbf{a}$ is the fast variable and $\mathbf{v},\mathbf{x}$ are the slow variables. In reverse time, the acceleration, $\mathbf{a}$ will be attracted to the slow manifold, given by
(101) $\mathbf{a}^{*}(\mathbf{x},\mathbf{v})=\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x})+\epsilon\frac{3}{2}\mathbf{v}\times\mathbf{B}(\mathbf{x})\times\mathbf{B}(\mathbf{x})+\mathcal{O}(\epsilon^{2}).$
@@
It is of interest to obtain forward-time trajectories of electrons along the slow manifold. A standard numerical integration of (99) will result in unstable trajectories on the fast time scale since any numerical error will push trajectories off of the slow-manifold, resulting in exponential growth of the acceleration. We instead apply FSNNs to learn the Fenichel normal form of the system in reverse time. We then obtain forward-time trajectories along the learned slow manifold by performing an integration of the slow variables along $y=0$ using (61). This test is therefore a more practical application of a data-driven closure.
We train an FSNN to learn the Abraham-Lorentz equations from data. The trajectory dataset, $\mathcal{D}$, consists of $10^{4}$ trajectories, corresponding to initial conditions randomly generated in the region, $\mathcal{R}$, specified by
@@
where $r=\sqrt{x_{1}^{2}+x_{2}^{2}}$ and $z=x_{3}$. $\epsilon$ is randomly generated log-uniformly in the range $[10^{-5},10^{-2}]$. The trajectories are integrated in reverse time using equations (100) using an adaptive RK45 solver set to an error tolerance of $10^{-8}$. The first 40 time steps batched in sequences of length 10 are used as fast-scale training data. These trajectories are also interpolated over a slower time scale with $\Delta t=0.025/\epsilon$ and $10^{4}$ sequences of length 10 are chosen at random for slow-scale training data. The manifold dataset, $\mathcal{M}$ consists of $10^{4}$ examples of data pairs $((\mathbf{x},\mathbf{v},\mathbf{a}),\epsilon=0)$ on the graph $\mathbf{a}=\mathbf{v}\times\mathbf{B}(\mathbf{x})$, where $\mathbf{x},\mathbf{v}$ are randomly generated in the region $\mathcal{R}$.
@@
(102) $B_{r}$ $=-\frac{z}{q(r,z)r}B_{0},\qquad B_{\phi}=\frac{R_{0}}{r}B_{0},\qquad B_{z}=\frac{(r-R_{0})}{q(r,z)r}B_{0},$
@@
(103) $\mathbf{B}=\frac{1}{R}\nabla\psi\times\mathbf{e}_{\phi}+\frac{F_{0}}{R}\mathbf{e}_{\phi}.$
@@
A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 50. The function $g$ on the right-hand-side of the neural ODE uses a single-layer feedforward neural network with hidden dimension 100 and is added to a bilinear form network with rank 1. The remaining feedforward neural networks on the right-hand-side of the neural ODE system each consist of 50 layers and the bilinear map has a rank of 4.
Figure 9 shows comparisons of the solutions in $r$-$z$ space between the trained network and trajectories computed using RK45 on the reverse-time system. Using the reverse-time trajectories, endpoints were supplied as initial conditions to the trained network to predict forward in time. For both the reverse-time predictions (red arrows) and forward-time predictions (green arrows), the trained model agrees well with the solution computed numerically. The corresponding $L_{2}$ errors are shown for position, velocity, and acceleration in Figure 10, demonstrating that errors are low when integrating on the slow time scale.
Table 2: Comparison of the true eigenvalues Jacobian of the fast dynamics to the eigenvalues of the negative Schur form of the trained model for the Abraham-Lorentz problem.
@@
Figure 11 shows comparisons of orbits between the trained model and orbits obtained when integrating the system by substituting the first two terms of the asymptotic expansion in (101). Note again that when integrating along the slow manifold, only the slow variable is evolved using the transformation $h$ and the closure $y=0$. The trained model is able to qualitatively capture the gyro-motion behavior while also accurately capturing the guiding center dynamics over the slow time scale. Despite training using trajectories over a relatively short time frame (2 or less gyration periods) due to the slow time scale instability of the reverse-time system, the trained network is highly capable of predicting long-time dynamics in forward-time for many more gyration cycles. We remind the reader that standard numerical methods cannot integrate the forward-time Abraham-Lorentz system along the slow manifold without experiencing an exponential growth instability. Direct integration of the asymptotic expansion removes the instability, but introduces an error dependent on $\epsilon$ due to the asymptotic approximation. This result highlights the unique capability of the proposed network.
@@
Figure 11: Comparison of orbits of the forward-time Abraham-Lorentz system (99) along the slow manifold for the first two terms of the asymptotic expansion in (101) (left) and the trained model (right) for two different initial conditions and choices of $\epsilon$.
@@ 6 Conclusions
In this paper, we introduced the fast-slow neural network (FSNN), which is a data-driven approach suitable for learning singularly perturbed dynamical systems where the fast-scale dynamics are dissipative. Our method enforces the existence of a trainable, attracting invariant slow manifold as a hard constraint. Invertible neural networks and neural ODEs are key components of the FSNN. We introduce bi-Lipschitz affine transformation (bLAT) layers for enforcing regularity of the invertible neural network. We also develop a negative Schur form network which parameterizes matrices with eigenvalues to the left of the imaginary axis. This in addition to the use of an additive L-stable diagonally-implicit Runge Kutta scheme for integration are key pieces in ensuring that the fast-scale dynamics are stable. The Schur form network also enables the use of back-substitution to efficiently handle the implicit solve. Analytical representation of the slow manifold enables efficient integration on the slow time scale. We demonstrate the FSNN on many examples that exhibit two timescales, including the Grad moment system from hydrodynamics, two-scale Lorenz96 equations for modeling atmospheric dynamics, and Abraham-Lorentz dynamics modeling radiation reaction of electrons in a magnetic field.
In order to effectively train the FSNN, a sufficient number of trajectories across multiple values of $\epsilon$ and initial conditions are necessary. Despite this, we demonstrate that the network can be used to discover a full-order model with long-term accuracy near a slow manifold when only short duration trajectories are available. Therefore the proposed network is also applicable to discover a reduced-order model as a unique closure discovery procedure.
In the examples we consider, the equations and slow manifold reduction in the limit $\epsilon=0$ are known a priori. Therefore, we were able to bypass an important shortcoming of our approach, which is the need to determine the dimensionality of the slow manifold. While a plethora of techniques exist in the literature [47], it would be interesting to combine them to our technique to tackle problems where the dimensionality is unknown.
A future direction of our work involves considering hyperbolic fast slow systems, which can be represented using a more general Fenichel normal form consisting of both stable and unstable directions.